OpenStudy (anonymous):

Find (in terms of constant a) lim (h -> 0) [(sqrt(7(a+h)) - (sqrt(7a)] / h

OpenStudy (phi):

there is a trick to this one. multiply top and bottom by \[ \sqrt{7(a+h)} + \sqrt{7a} \] we will be using (a-b)(a-b)= a^2 - b^2 to simplify the top

OpenStudy (phi):

**we will be using (a-b)(a+b)= a^2 - b^2 to simplify the top